# NIMO: a Nonlinear Interpretable MOdel

### v2

@@ 1 Introduction
Over the past decade, neural networks have achieved remarkable success across domains, from computer vision (Krizhevsky et al., 2012) to natural language processing (Vaswani et al., 2017). At the same time, their deployment in high-stakes domains such as healthcare has created a pressing demand for interpretability (Esteva et al., 2017). However, neural networks in general lack interpretability of the predictions in terms of the input features, hence the term “black-box” models. In contrast, classical models such as linear regression and decision trees are often considered as highly interpretable because of their transparent structure and simple decision rules, but they can lack predictive power on complex, high-dimensional, or highly nonlinear tasks (Hastie, 2009; Breiman et al., 2017). This tension between accuracy and interpretability has motivated much recent work on models and methods that combine the expressive power of neural networks with the transparency of simpler models (Lemhadri et al., 2021; Thompson et al., 2023; Wu et al., 2017; 2021). Meanwhile, many model-agnostic, post-hoc explainers have been proposed, such as SHAP (Lundberg & Lee, 2017) and LIME (Ribeiro et al., 2016), providing instance-level feature attributions for arbitrary predictors. But these post-hoc explanations are approximations and may depend on the choices of parameters, such as the background distribution and kernel width. Among various ways to achieve interpretability, feature effects (Scholbeck et al., 2019) provide a direct, quantitative description of how each feature influences the prediction, something post-hoc methods can approximate but do not guarantee.
@@
Since the network parameters $\bm{u}$ and the linear coefficients $\bm{\beta}$ are tightly coupled, jointly optimizing them is non-trivial. To address this, we develop an optimization method based on parameter elimination. Specifically, we derive a closed-form expression for the linear coefficients $\bm{\beta}$ as a function of the NN parameters $\bm{u}$. Substituting this expression back into the objective reduces the problem to optimizing only over the NN parameters. To further enhance interpretability, we impose sparsity by formulating the problem as adaptive ridge regression (Grandvalet, 1998). In the network, we further apply group $\ell_{2}$ regularization (Yuan & Lin, 2006) to the weight matrix of the first fully connected layer, encouraging feature-level sparsity and clarifying how inputs contribute to predictions.
@@ 2 Related Work
Interpretability in machine learning. Interpretability refers to the capacity to express what a model has learned and the factors influencing its outputs in a manner that is clear and understandable to humans. In the literature, the two terms interpretability and explainability are sometimes used interchangeably. However, they have subtle yet important differences. Explainability focuses on providing reasons for the model’s output. Most of the methods, especially in deep learning, provide post-hoc explanations. The most popular ones include LIME (Ribeiro et al., 2016), SHAP (Lundberg & Lee, 2017) and Grad-CAM (Selvaraju et al., 2020). On the other hand, interpretability focuses on understanding the inner mechanisms and decision processes of the model, and most of the methods that claimed to be interpretable involve inherently understandable models, such as decision trees (Wu et al., 2017; 2021), Lasso (Tibshirani, 1996) and Explainable Boosting Machines (Lou et al., 2013). The literature on interpretable machine learning is vast and we refer readers to Rudin et al. (2022) and Molnar (2025) for a more in-depth review.
@@
Hybrid models. Several works have proposed combining linear models with neural networks to improve interpretability. Neural Additive Models (NAMs) (Agarwal et al., 2021) extend generalized additive models by learning one neural network per feature, whose outputs are summed to form the prediction. This design preserves interpretability at the feature level but does not have the ability to capture feature interactions. LassoNet (Lemhadri et al., 2021) integrates a neural network with Lasso regression by enforcing that a feature can be used in the nonlinear part only if its corresponding linear coefficient is nonzero. Interpretable Mesomorphic Networks (IMN) (Kadra et al., 2024) use a hypernetwork to predict linear coefficients on a per-instance basis. While this increases flexibility, it sacrifices global interpretability of the coefficients, making it difficult to recover clear baseline effects. In contrast, our proposed model (NIMO) preserves the global interpretability of linear coefficients while augmenting them with nonlinear corrections. This unifies global summaries via MEM with instance-level explanations within a single coherent framework.
@@ 3.3 Training via parameter elimination
Consider the data $X\in\mathbb{R}^{n\times d}$ and targets $\mathbf{y}\in\mathbb{R}^{n}$. As argued before, we can model the $d$ networks with one network $g_{\bm{u}}$ and represent its outputs in a matrix $G_{\bm{u}}=g_{\bm{u}}(X)\in\mathbb{R}^{n\times d}$. The model in Eq. 2 can then be re-written in matrix form as $f(X)=B_{\bm{u}}\bm{\beta}$, where $B_{\bm{u}}=X+X\circ G_{\bm{u}}$ and $\circ$” denotes element-wise multiplication. Let us first consider the standard ridge regression setting: $\mathcal{L}(\bm{\beta},\bm{u})=\lVert\mathbf{y}-B_{\bm{u}}\bm{\beta}\rVert^{2}+\lambda\lVert\bm{\beta}\rVert^{2}$. To separate the optimization over $\bm{\beta}$ and $\bm{u}$, we take inspiration from the profile likelihood approach (Venzon & Moolgavkar, 1988; Murphy & Van der Vaart, 2000). The key idea is to eliminate $\bm{\beta}$ by first solving for its closed-form expression in terms of $\bm{u}$, and then substituting it back into the objective. This reduces the problem to an optimization over $\bm{u}$ only:
@@
Unlike ridge regression, Lasso does not admit a closed-form solution for $\bm{\beta}$. To eliminate $\bm{\beta}$ in this setting, we use adaptive ridge regression, which admits at each step a closed-form expression for $\bm{\beta}$ in terms of $\bm{u}$ and is equivalent to Lasso at the optimum (Grandvalet, 1998); see Appendix A for a detailed proof. This enables us to efficiently optimize the problem using gradient descent while still enforcing sparsity. The full procedure is summarized in Algorithm 1 in the Appendix. A well-known limitation of Lasso is the over-shrinkage effect, where large coefficients are excessively penalized (Fan & Li, 2001). In some applications, this can lead to significant bias and degrade model performance. To mitigate this issue, we can extend the $\ell_{1}$ norm regularization to a sub-$\ell_{1}$ pseudo-norm, which penalizes large coefficients less aggressively while still promoting sparsity. A detailed introduction to this approach is provided in Appendix B.3.
@@
The same idea can be extended to any generalized linear model of the form: $f(\mathbb{E}[Y|\mathbf{x}])=\beta_{0}+\beta_{1}x_{1}+\dots+\beta_{p}x_{p}=\mathbf{x}^{T}\bm{\beta}$, where $f$ is a link function. For simplicity, consider logistic regression with $\ell_{1}$ penalty. As before, we first get a closed-form expression for $\bm{\beta}$ in terms of $\bm{u}$ and then substitute it back to optimize over the neural network parameters $\bm{u}$. What is different is that in each optimization step we use the iteratively reweighted least squares (IRLS) surrogate (Green, 1984; McCullagh & Nelder, 1989), which is a weighted least-squares approximation of the original problem:
@@ Methods:
In this paper, we compare our method against Lasso (Tibshirani, 1996), a vanilla neural network (NN(Goodfellow et al., 2016), and several state of the art interpretable approaches, namely LassoNet (Lemhadri et al., 2021), Neural Additive Models (NAMs(Agarwal et al., 2021) and Interpretable Mesomorphic Networks (IMNs(Kadra et al., 2024). LassoNet encourages the network to use only a subset of the available input features by enforcing sparsity with a novel objective. NAMs learn a linear combination of neural networks that each attend to a single input feature. IMNs optimize deep hypernetworks to generate explainable linear models on a per-instance basis.
@@ 4.2 Hypotheses and experiment results
We first examine the predictive performance of NIMO. The box plot of the mean squared error for each method is shown in Figure 5. NIMO generally achieves performance compatible with the other methods and attains the best performance on the superconductivity dataset. To assess interpretability, we first visualize the marginal effects at the mean (MEM) obtained from Lasso and NIMO for the diabetes and Boston housing datasets in Figure 6. Recall that MEM estimates marginal effects at the average values of the input features, reflecting the additive population-level contribution of each feature. The visualizations reveal that the MEM from NIMO differ substantially from those of Lasso, reflecting NIMO’s ability to capture nonlinear feature interactions that Lasso cannot represent. For the superconductivity dataset, which involves a much larger set of features, the corresponding MEM visualization is provided in the Appendix.
We then use MEM to rank the features on the diabetes and Boston housing datasets. For comparison, we also apply the post-hoc explainer SHAP (Lundberg & Lee, 2017) to our model and rank features according to the global SHAP values. The results are shown in Table 2. On the diabetes dataset, all methods behave similarly in identifying the most important features, which is expected since the data lacks complex nonlinear interactions. This is further supported by the comparable predictive performance across methods, as shown in Figure 5 (left). Notably, SHAP produces identical feature rankings to MEM for NIMO. In contrast, for the Boston housing dataset, the methods yield inconsistent feature rankings, suggesting that nonlinear interactions play a crucial role in explaining the data. Although MEM can also be computed for other nonlinear methods, they do not provide the same level of intelligible interpretability as our model. For NIMO, SHAP rankings are largely consistent with MEM, though not perfectly aligned. The most pronounced difference appears with LassoNet, which assigns high importance to features considered least relevant by other methods, such as INDUS. We attribute this to how LassoNet handles nonlinearities.
@@ 5 Conclusions
In this paper, we propose a nonlinear interpretable model (NIMO). Unlike other hybrid approaches, our model preserves the interpretability of linear coefficients, capturing the additive contribution of each feature to the prediction, while also accommodating nonlinear interactions. Interpretability is formalized through marginal effects, with the marginal effects at the mean (MEM) providing global, population-level summaries that align directly with the linear coefficients. To enable effective training, we introduce an optimization algorithm based on parameter elimination, and enforce sparsity through adaptive ridge regularization. We demonstrate the effectiveness of the proposed approach on both synthetic and real datasets, showing competitive predictive performance while delivering meaningful and globally consistent interpretations. A key limitation of the approach is that it assumes the data can be described as a linear model with nonlinearity corrections; highly nonlinear datasets or features with self-interactions (e.g., $x_{j}f(x_{j})$) may not be fully captured. In such cases, careful feature engineering may be required to preserve interpretability.
@@ Appendix A Equivalence of Adaptive Ridge regression and Lasso
The equivalence in the solution of Adaptive ridge regression and Lasso regression is a well-known fact (Tibshirani, 1996; Grandvalet, 1998; Grandvalet & Canu, 1998). Since both our model relies on this fact both in the regression and classification settings, we report its proof below.
@@ B.1 Regression setting: adaptive ridge regression
where $\nu_{i}>0$ are the penalty parameters for each coefficient $\beta_{i}$, and $\lambda>0$ is a predefined parameter. To avoid divergent solutions we use the re-parameterization employed in Grandvalet (1998); Grandvalet & Canu (1998):
@@ C.2 Practical Implementation
We implement our model with PyTorch Lightning (Falcon & The PyTorch Lightning team, 2019). While the model definition is simple, in practice, several strategies are needed to make the training more efficient and effective.
@@ D.1 Methods implementation
We implement the naive neural network (NN) with PyTorch Lightning (Falcon & The PyTorch Lightning team, 2019). For a fair comparison, the NN has the same structure as the nonlinear part in our NIMO model, except no positional encoding. We use Adam (Kingma, 2014) as the default optimizer. The learning rate is 1e-3 for synthetic experiments, and 5e-3 for real-world datasets. Strong dropout regularization (p=0.6) is used to avoid overfitting.

### v3

@@ 1 Introduction
Over the past decade, neural networks have achieved remarkable success across domains, from computer vision (Krizhevsky et al., 2012) to natural language processing (Vaswani et al., 2017). At the same time, their deployment in high-stakes domains such as healthcare has created a pressing demand for interpretability (Esteva et al., 2017). However, neural networks in general lack interpretability of the predictions in terms of the input features, hence the term “black-box” models. In contrast, classical models such as linear regression and decision trees are often considered as highly interpretable because of their transparent structure and simple decision rules, but they can lack predictive power on complex, high-dimensional, or highly nonlinear tasks (Hastie et al., 2009; Breiman et al., 2017). This tension between accuracy and interpretability has motivated much recent work on models and methods that combine the expressive power of neural networks with the transparency of simpler models (Lemhadri et al., 2021; Thompson et al., 2023; Wu et al., 2017; 2021). Meanwhile, many model-agnostic, post-hoc explainers have been proposed, such as SHAP (Lundberg and Lee, 2017) and LIME (Ribeiro et al., 2016), providing instance-level feature attributions for arbitrary predictors. But these post-hoc explanations are approximations and may depend on the choices of parameters, such as the background distribution and kernel width. Among various ways to achieve interpretability, feature effects (Scholbeck et al., 2019) provide a direct, quantitative description of how each feature influences the prediction, something post-hoc methods can approximate but do not guarantee.
@@
Since the network parameters $\bm{u}$ and the linear coefficients $\bm{\beta}$ are tightly coupled, jointly optimizing them is non-trivial. To address this, we develop an optimization method based on parameter elimination. Specifically, we derive a closed-form expression for the linear coefficients $\bm{\beta}$ as a function of the NN parameters $\bm{u}$. Substituting this expression back into the objective reduces the problem to optimizing only over the NN parameters. To further enhance interpretability, we impose sparsity by formulating the problem as adaptive ridge regression (Grandvalet, 1998). In the network, we further apply group $\ell_{2}$ regularization (Yuan and Lin, 2006) to the weight matrix of the first fully connected layer, encouraging feature-level sparsity and clarifying how inputs contribute to predictions.
@@ 2 Related Work
Interpretability in machine learning. Interpretability refers to the capacity to express what a model has learned and the factors influencing its outputs in a manner that is clear and understandable to humans. In the literature, the two terms interpretability and explainability are sometimes used interchangeably. However, they have subtle yet important differences. Explainability focuses on providing reasons for the model’s output. Most of the methods, especially in deep learning, provide post-hoc explanations. The most popular ones include LIME (Ribeiro et al., 2016), SHAP (Lundberg and Lee, 2017) and Grad-CAM (Selvaraju et al., 2020). On the other hand, interpretability focuses on understanding the inner mechanisms and decision processes of the model, and most of the methods that claimed to be interpretable involve inherently understandable models, such as decision trees (Wu et al., 2017; 2021), Lasso (Tibshirani, 1996) and Explainable Boosting Machines (Lou et al., 2013). The literature on interpretable machine learning is vast and we refer readers to Rudin et al. (2022) and Molnar (2025) for a more in-depth review.
@@
Hybrid models. Several works have proposed combining linear models with neural networks to improve interpretability. Neural Additive Models (NAMs) (Agarwal et al., 2021) extend generalized additive models (GAMs) by learning one neural network per feature, whose outputs are summed to form the prediction. This design preserves interpretability at the feature level but does not have the ability to capture feature interactions. NODE-GAM (Chang et al., 2021) is also a neural generalized additive model that uses differentiable tree ensembles to learn interpretable univariate and pairwise shape functions. However, similar to NAM, both of their interpretability comes from the shape function, which is input dependent, and cannot provide a global population-level summary. LassoNet (Lemhadri et al., 2021) integrates a neural network with Lasso regression by enforcing that a feature can be used in the nonlinear part only if its corresponding linear coefficient is nonzero. Interpretable Mesomorphic Networks (IMN) (Kadra et al., 2024) use a hypernetwork to predict linear coefficients on a per-instance basis. While this increases flexibility, it sacrifices global interpretability of the coefficients, making it difficult to recover clear baseline effects. Interpretable Mixture of Experts (IME) (Ismail et al., 2022) routes each sample to an interpretable expert (e.g., a linear model), ensuring that for that sample the entire decision process is an exact and transparent explanation of which expert was chosen and why. While IME can maintain high accuracy and provide faithful local explanations, its global interpretability remains limited. In contrast, our proposed model (NIMO) preserves the global interpretability of linear coefficients while augmenting them with nonlinear corrections. This unifies global summaries via MEM with instance-level explanations within a single coherent framework.
@@ 3.2 Interpretability via marginal effects
While MEM is widely used as a global effect measure in statistics, it is important to acknowledge its general limitations, as discussed in Scholbeck et al. (2024). MEM is meaningful and interpretable in models that include an explicit global linear component, but for highly nonlinear functions the MEM may fail to reflect the true relevance of individual features. However, NIMO is explicitly designed to avoid such cases: its architecture explicitly separates a global linear component (captured by $\beta$) from local, per-instance nonlinear adjustments. This structural constraint ensures that the global linear contribution remains identifiable and stable. Consequently, MEM remains appropriate and interpretable within NIMO’s model class, even though it may be unsuitable as a universal feature-importance metric for fully unconstrained nonlinear functions.
@@ 3.3 Training via parameter elimination
Consider the data $X\in\mathbb{R}^{n\times d}$ and targets $\mathbf{y}\in\mathbb{R}^{n}$. As argued before, we can model the $d$ networks with one network $g_{\bm{u}}$ and represent its outputs in a matrix $G_{\bm{u}}=g_{\bm{u}}(X)\in\mathbb{R}^{n\times d}$. The model in Eq. 2 can then be re-written in matrix form as $f(X)=B_{\bm{u}}\bm{\beta}$, where $B_{\bm{u}}=X+X\circ G_{\bm{u}}$ and $\circ$” denotes element-wise multiplication. Let us first consider the standard ridge regression setting: $\mathcal{L}(\bm{\beta},\bm{u})=\lVert\mathbf{y}-B_{\bm{u}}\bm{\beta}\rVert^{2}+\lambda\lVert\bm{\beta}\rVert^{2}$. To separate the optimization over $\bm{\beta}$ and $\bm{u}$, we take inspiration from the profile likelihood approach (Venzon and Moolgavkar, 1988; Murphy and Van der Vaart, 2000). The key idea is to eliminate $\bm{\beta}$ by first solving for its closed-form expression in terms of $\bm{u}$, and then substituting it back into the objective. This reduces the problem to an optimization over $\bm{u}$ only:
@@
Unlike ridge regression, Lasso does not admit a closed-form solution for $\bm{\beta}$. To eliminate $\bm{\beta}$ in this setting, we use adaptive ridge regression, which admits at each step a closed-form expression for $\bm{\beta}$ in terms of $\bm{u}$ and is equivalent to Lasso at the optimum (Grandvalet, 1998); see Appendix A for a detailed proof. This enables us to efficiently optimize the problem using gradient descent while still enforcing sparsity. The full procedure is summarized in Algorithm 1 in the Appendix. A well-known limitation of Lasso is the over-shrinkage effect, where large coefficients are excessively penalized (Fan and Li, 2001). In some applications, this can lead to significant bias and degrade model performance. To mitigate this issue, we can extend the $\ell_{1}$ norm regularization to a sub-$\ell_{1}$ pseudo-norm, which penalizes large coefficients less aggressively while still promoting sparsity. A detailed introduction to this approach is provided in Appendix B.3.
@@
The same idea can be extended to any generalized linear model of the form: $f(\mathbb{E}[Y|\mathbf{x}])=\beta_{0}+\beta_{1}x_{1}+\dots+\beta_{p}x_{p}=\mathbf{x}^{T}\bm{\beta}$, where $f$ is a link function. For simplicity, consider logistic regression with $\ell_{1}$ penalty. As before, we first get a closed-form expression for $\bm{\beta}$ in terms of $\bm{u}$ and then substitute it back to optimize over the neural network parameters $\bm{u}$. What is different is that in each optimization step we use the iteratively reweighted least squares (IRLS) surrogate (Green, 1984; McCullagh and Nelder, 1989), which is a weighted least-squares approximation of the original problem:
@@ Methods:
In this paper, we compare our method against Lasso (Tibshirani, 1996), a vanilla neural network (NN(Goodfellow et al., 2016), and several state of the art interpretable approaches, namely LassoNet (Lemhadri et al., 2021), Neural Additive Models (NAMs(Agarwal et al., 2021) and Interpretable Mesomorphic Networks (IMNs(Kadra et al., 2024).
@@
LassoNet encourages the network to use only a subset of the available input features by enforcing sparsity with a novel objective. NAMs learn a linear combination of neural networks that each attend to a single input feature. IMNs optimize deep hypernetworks to generate explainable linear models on a per-instance basis.
@@ 4.2 Hypotheses and experiment results
We first examine the predictive performance of NIMO. The box plot of the mean squared error for each method is shown in Figure 5. NIMO generally achieves performance compatible with the other methods and attains the best performance on the superconductivity dataset. To assess interpretability, we first visualize the marginal effects at the mean (MEM) obtained from Lasso and NIMO for the diabetes and Boston housing datasets in Figure 6. Recall that MEM estimates marginal effects at the average values of the input features, reflecting the additive population-level contribution of each feature. We observe that the MEM profiles of Lasso and NIMO differ across several features. These differences do not necessarily imply the presence of nonlinear feature interactions; rather, both datasets contain groups of strongly correlated predictors, making feature selection inherently unstable. Consequently, Lasso and NIMO may highlight different features within the same correlated group, even if their predictive roles are similar. In Appendix D.8, we experimentally showcase the degree of nonlinearities in each dataset. For the superconductivity dataset, which involves a much larger set of features, the corresponding MEM visualization is provided in the Appendix in Figure 14.
We then use MEM to rank the features on the diabetes and Boston housing datasets. For comparison, we also apply the post-hoc explainer SHAP (Lundberg and Lee, 2017) to our model and rank features according to the global SHAP values. The results are shown in Table 2. On the diabetes dataset, all methods behave similarly in identifying the most important features, which is expected since the data lacks complex nonlinear interactions. This is further supported by the comparable predictive performance across methods, as shown in Figure 5 (left). Notably, SHAP produces identical feature rankings to MEM for NIMO. In contrast, for the Boston housing dataset, the methods yield inconsistent feature rankings, suggesting that nonlinear interactions play a crucial role in explaining the data. Although MEM can also be computed for other nonlinear methods, they do not provide the same level of intelligible interpretability as our model. For NIMO, SHAP rankings are largely consistent with MEM, though not perfectly aligned. The most pronounced difference appears with LassoNet, which assigns high importance to features considered least relevant by other methods, such as INDUS. We attribute this to how LassoNet handles nonlinearities.
@@ 4.3 Ablation study
4.4 Limitations
The proposed approach is designed to preserve the global, population-level interpretability inherited from linear models while simultaneously learning nonlinear corrections to linear predictions through neural networks, thereby enhancing the expressive capacity of linear models. Thus, the underlying assumption is that the data can be well described by adding nonlinear refinements on top of a linear model. In settings dominated by strong self-interactions or highly nonlinear regimes, NIMO cannot fully capture the underlying relationships in an interpretable manner. Assume the data-generating process contains a self-interacting term of the form $x_{j}f(x_{j})$, where $f$ is a linear or nonlinear function (e.g. $x\sin(x)$ or $x^{2}$). To maintain interpretability of the linear coefficients, we imposed the constraint that $g_{\bm{u}_{j}}(\mathbf{\mathbf{x}_{-j}})$ does not depend on $x_{j}$, which prevents NIMO from learning such self-interacting nonlinearities.
However, if such a feature is known to play an important role in the dataset, simple feature engineering can resolve the issue. To experimentally and explicitly illustrate both the limitation of NIMO and a straightforward remedy, we conduct an ablation study on a synthetic dataset. The data are generated as
$y=x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+\epsilon.$ (9)
Empirically, we demonstrate that naively applying NIMO to this dataset results in poor performance. However, after applying a polynomial basis expansion of degree 2, NIMO can perfectly solve the problem while still providing global, population-level interpretability through the linear coefficients $\bm{\beta}$. More detailed experimental results can be found in Appendix D.6.
@@ 5 Conclusions
In this paper, we propose a nonlinear interpretable model (NIMO). Unlike other hybrid approaches, our model preserves the interpretability of linear coefficients, capturing the additive contribution of each feature to the prediction, while also accommodating nonlinear interactions. Interpretability is formalized through marginal effects, with the marginal effects at the mean (MEM) providing global, population-level summaries that align directly with the linear coefficients. To enable effective training, we introduce an optimization algorithm based on parameter elimination, and enforce sparsity through adaptive ridge regularization. We demonstrate the effectiveness of the proposed approach on both synthetic and real datasets, showing competitive predictive performance while delivering meaningful and globally consistent interpretations.
@@ Appendix A Equivalence of Adaptive Ridge regression and Lasso
The equivalence in the solution of Adaptive ridge regression and Lasso regression is a well-known fact (Tibshirani, 1996; Grandvalet, 1998; Grandvalet and Canu, 1998). Since both our model relies on this fact both in the regression and classification settings, we report its proof below.
@@ B.1 Regression setting: adaptive ridge regression
where $\nu_{i}>0$ are the penalty parameters for each coefficient $\beta_{i}$, and $\lambda>0$ is a predefined parameter. To avoid divergent solutions we use the re-parameterization employed in Grandvalet (1998); Grandvalet and Canu (1998):
@@ C.2 Practical Implementation
We implement our model with PyTorch Lightning (Falcon and The PyTorch Lightning team, 2019). While the model definition is simple, in practice, several strategies are needed to make the training more efficient and effective.
@@ D.1 Methods implementation
We implement the naive neural network (NN) with PyTorch Lightning (Falcon and The PyTorch Lightning team, 2019). For a fair comparison, the NN has the same structure as the nonlinear part in our NIMO model, except no positional encoding. We use Adam (Kingma, 2014) as the default optimizer. The learning rate is 1e-3 for synthetic experiments, and 5e-3 for real-world datasets. Strong dropout regularization (p=0.6) is used to avoid overfitting for small scale datasets (including synthetic datasets, diabetes and Bostong housing datasets). For the relatively large superconductivity dataset, small dropout rate (p=0.1) is used.